# Dust destruction by the supernova remnant forward shock in a turbulent interstellar medium

### v1

@@ 3.4 Summary
To summarize, the SNR forward shock dust destruction depends on the properties of the surrounding medium (ISM density and turbulent Mach number), the dust grain species, and the considered destruction mechanisms. If only one of these parameters is varied, the dust destruction efficiency can change significantly. Generally, we find that the dust destruction efficiency decreases with lower ISM densities and higher turbulent Mach numbers. It is further reduced if carbonaceous instead of silicate dust is considered. Moreover, grain-grain collisions always generate many grains with sizes lower than 5 nm and can have a large impact on the destruction rate in intermediate and high density environments (up to 36%).
The lowest forward shock dust destruction in the first 10 kyr is found in the low density ($n=1\,\mathrm{cm^{-3}}$) and strongly turbulent ($\mathcal{M}=3$) environment, where only 27% of carbonaceous dust is destroyed. With 0.85 M of carbonaceous dust, the least dust mass is being destroyed in the $n=1\,\mathrm{cm^{-3}}$ and $\mathcal{M}=0.3$ case. However, the forward shock still has a high velocity of $v>250\,\mathrm{km\,s^{-1}}$ and is still hot at the end of these simulations, so that especially kinetic and thermal sputtering will process much more dust beyond the considered period. Compared to the average dust mass of 0.1–1 M that is observed to be produced by SNRs and that still needs to survive the high velocity reverse shock, we conclude that SNRs are rather net dust destroyers, in particular if higher gas densities are considered.
@@ 4.1 Previous studies
Table 2: Comparison of this study to different forward shock dust destruction studies after 10 kyr.444[1] Dedikov2025, [2] Kirchschlager2024a, [3] Kirchschlager2022, [4] Martinez-Gonzalez2019. Compared are the ISM inhomogeneity model, the average ISM density ($n$), the resolution, whether grain-grain collision are considered or not ($g$), the initial grain-size distribution (GSD), and the destroyed dust mass (and rate) after the first 10 kyr of SNR evolution for $n\approx 1\,\mathrm{cm^{-3}}$ cases.
@@
Such a study has been performed by Dedikov2025, who use a similar approach to ours by simulating the forward shock dust destruction over 100 kyr in a clumpy instead of turbulent medium, which is generated by fractal cubes. They use different inhomogeneity strengths and their densest clumps have roughly the same density contrasts between high and low density regions as our supersonic Mach number simulations. They do not consider magnetic fields nor grain-grain collisions, only kinetic and thermal sputtering. In their paper, they conclude that dense clumps can shield the dust efficiently, resulting in up to $\sim$20% less destruction over 100 kyr for an ISM with $n=1\,\mathrm{cm^{-3}}$ on average. After 10 kyr, they find 1.17 M ($\sim$39%) of destroyed dust when considering a homogeneous medium, decreasing with increasing inhomogeneity down to 0.67 M ($\sim$32%) of dust with the densest clumps so that these clumps already shield $\sim$7% of the dust at 10 kyr. Their considered grain species is not mentioned but the results match reasonably well to our carbonaceous (0.79–0.91 M or 25–33%) grain simulations with $n=1\,\mathrm{cm^{-3}}$ that include only sputtering, for which we also find significant filament shielding of 8% at 10 kyr. In an intermediate density environment ($n=10\,\mathrm{cm^{-3}}$) with rather dense ISM clumps, $\sim$4.5 M ($\sim$50%) of dust is destroyed in the first 10 kyr, also fitting reasonably well to our results of the intermediate density case with strong turbulence and carbonaceous grains (3.5–4.2 M or 46–63% of dust destroyed). This supports the conclusion that generally more dust is being destroyed at higher ISM densities. At 100 kyr, however, this trend can reverse according to their study where 8.0–11.4 M dust is destroyed in the low density, but only 7.5 M in the inhomogeneous intermediate density case.
@@
If our sputtering results for carbonaceous grains would also align quite well with the results of Dedikov2025 beyond 10 kyr, over 8 M of dust would be destroyed at 100 kyr, which is almost an order of magnitude more than what SNRs are observed to produce. Nevertheless, we note that the filaments in our study could behave differently than the clumpy medium of Dedikov2025 at lower shock velocities, and that we cannot estimate the dust evolution in our own simulations that include grain-grain collisions beyond 10 kyr based on this comparison, as this destruction mechanism could become more efficient at later SNR stages (see for example our $n=100\,\mathrm{cm^{-3}}$ case).
@@
Kirchschlager2024a also investigated single SNRs in a turbulent ISM, similarly to us. However, they used SN-driven turbulence, which brings the caveats mentioned in Sect. 1 with it. After turbulence driving, they inject an SN in an either low or moderate density area of $n=0.03\text{ or }0.7$ cm-3, respectively. Thus, although their average ISM density is always lower than our lowest density case ($n=1$ cm-3), the SN injection in their moderately dense region might be comparable to our simulations because their SNR does not encounter many voids whereas we set off the SN in a void. Their dust properties and considered dust destruction mechanisms are similar to our study, but they only consider silicate grains.
@@
After 10 kyr, they find a destroyed dust mass of 2.8 M in the moderate density simulation, which is almost 2 times more than the destroyed silicate grain mass in our high Mach turbulence study (1.5 M). This disparity can come from multiple factors. Most importantly, the turbulence driving is different so that their produced filaments seem to be much wider than ours. Furthermore, the average ISM density in the box is lower with the SN set off in a higher density environment. Thus, the dust destruction efficiency can change significantly (Kirchschlager2024a) and the encountered dust mass could be different. Moreover, their simulation is not fine-tuned to the first 10 kyr, meaning that the temporal and spatial resolution is rather coarse (250 yr and 0.5 pc) compared to our study (logarithmic time steps between 0.15–120 yr depending on the evolution time and 0.044 pc in the $n=1$ cm-3 cases), which can cause inaccuracies of the gas drag. With a lower spatial resolution and thus bigger grid cells, the dust velocity in a single cell is averaged over more grains, resulting in larger deviations compared to the real grain velocities. If the temporal resolution is too low, grains that are dragged with the forward shock may not be as efficiently decelerated by encountered ISM filaments as the gas so that the grains could be able to overtake the shock. This would leave the high-velocity grains exposed in a rather stationary ISM, leading to high relative gas-grain and grain-grain velocities.
@@
Thus, the higher destruction efficiency of Kirchschlager2024a could be explained by an insufficient resolution. Other differences can add to these inaccuracies. First of all, the extrapolation from the 2D to the 3D dust mass is different and induces uncertainties. We multiply the averaged destruction efficiency of the three investigated 2D planes with the 3D SNR-encountered dust mass, whereas Kirchschlager2024a fully extrapolate the dust mass from one 2D plane, using a correction factor they derived from comparing the SNRs in a homogeneous and an inhomogeneous medium. Thus, the real 3D destruction could be different. Secondly, they do not include partial vaporization, which was implemented in Paperboats by Kirchschlager2023. This reduces the dust destruction for the largest grains in our simulations slightly. Thirdly, the penetration depth of gas ions into dust grains was calculated considering oxygen ions in their study, but updated in our study to consider hydrogen instead of oxygen ions, which have a different penetration depth. This value alters the size-dependent sputtering, so that significantly more small silicate grains are destroyed. Lastly, it should be mentioned that the dust destruction in a turbulent ISM also depends on the detailed ISM environment, as we have shown with the three different orthogonal planes in this paper. The differences in the destruction efficiencies could therefore also partly come from the different surrounding environment and position of the filaments.
@@
A similarity to our study, however, is that the derived dust destruction efficiencies of Kirchschlager2024a have not plateaued yet after 10 kyr, but only start to plateau after 50 kyr in their study. Nevertheless, they find that some dust will keep being destroyed even when their simulations finish at 1 Myr, where a total of 37 M of dust is destroyed, which is $\sim$13.5 times more than what they find at 10 kyr. If this trend is applied to our high Mach number and low density ISM silicate simulation, it would mean that even in one of the lowest dust destruction cases, the SNR would destroy $1.5\,\mathrm{M_{\odot}}\times 13.5\approx 20\,\mathrm{M_{\odot}}$ of dust after 1 Myr, being 20 times higher than the observed dust production in SNRs. This implies, once again, that SNRs are rather net dust destroyers.
@@
Another useful comparison can be made to works that consider pre-SN feedback. Martinez-Gonzalez2019 let a wind-driven bubble evolve in a homogeneous ISM with a density of 1 cm-3 before injecting the SN. The bubble’s shocked region is compressed and has a density that is roughly 3 to 4 orders of magnitude higher than the interior of the bubble in which the SN explodes, resulting in roughly the same difference between filaments and voids as in our low density and supersonic turbulence simulations. It expands up to 23 pc in radius, meaning that it sweeps up $\sim$12.6 M of dust during the stellar wind evolution. They find that $\sim$0.36 M of dust (2.9% of the shell’s dust mass) is destroyed in the first 10 kyr after SN explosion if only thermal sputtering is considered and a mixture of silicate and carbonaceous grains is used. Although we also find that dust can be shielded by filaments, significantly more dust is shielded by their stellar wind shell than in our sputtering-only results with high Mach turbulence for either grain species (0.88–1.5 M and 25–45%). However, their bubble is larger and its shell is with roughly 3–5 pc much thicker than our filaments, meaning that it can slow down the shock more efficiently. As a result, most of the dust in the shell has not yet been encountered after 10 kyr so that this dust cannot be destroyed. Furthermore, kinetic sputtering is not included by them, which could be efficient and increase the destroyed dust fraction when the SNR forward shock starts to encounter the wind-driven bubble. At this time point, grain-grain collisions could also become efficient due to the high dust densities and velocities, which would further increase the dust destruction.
Other studies that investigate the dust destruction of SNR forward shocks in a slightly inhomogeneous or turbulent ISM are Vasiliev2024 and Hu2019, respectively. The former uses the same code as Dedikov2025 but only introduces small density perturbations of 10% and does not compare their results to a homogeneous case so that the filament shielding efficiency cannot be assessed. Since they only study a monosized dust population, we cannot easily compare their study with ours because we use an MRN-distributed dust population, which leads to significantly different dust behavior. Hu2019 drive their turbulence with SNe, similar to Kirchschlager2024a. However, they study the lifetime of silicate and carbonaceous dust grains over a total of 400 Myr and several SNe. Thus, their found destruction efficiencies are also not easily comparable to our study that only covers the first 10 kyr. However, similarly to us, they find that silicates are better destroyed than carbonaceous grains.
@@ 4.2 Caveats
  • The pre-SN effects of stellar winds and radiation are only mimicked simplistically in our study by setting off the SNe in voids because a full pre-SN feedback model is out of the scope of this paper. Especially the weakly turbulent and homogeneous runs do not incorporate realistic pre-SN structures. For a forthcoming model, we want to improve this by allowing stellar winds and radiation to evolve naturally in separate simulations, and letting the SN explode inside of the wind-driven bubble.
@@
  • Although we consistently set off all SNe in voids, the detailed surrounding ISM structures are not the same, since the turbulence-driving results in different filamentary shapes for each simulation. This affects the 3D SNR evolution in the first 10 kyr and therefore also the dust destruction, resulting in an uncertainty, which could be quantified best by studying many different 3D ISM environments for every parameter set. Since the needed resources for this would be too high, we assess this uncertainty with our investigation of the three orthogonal 2D planes, as discussed in Appendix D. We find a maximum uncertainty of 9.4% and include the standard deviation between those planes in our dust destruction results of Tables 1 and 4. This implies that the SNR dust destruction cannot be precisely predicted based on the initial parameters like ISM density and turbulent Mach number, but needs to take the exact SN ejection location (see also Appendix B) and the surrounding environment structures into account.
@@
  • Our post-processing approach does not allow the dust to feed heavy elements coming from destroyed dust grains back to the gas, which physically should happen and induce for example a faster cooling of the gas. The exact impact of the dust feedback is so far unknown but might be significant. Since on-the-fly approaches are more expensive than post-processing simulations, we use the latter for our calculations to include a more extensive dust destruction model. However, this brings the drawbacks of no dust-to-gas feedback and coarser time steps, which gives inaccuracies in the destruction, transport, and charge processes that we try to minimize by taking as many logarithmically spaced snapshots as computationally feasible. A comprehensive on-the-fly approach is currently being implemented, tested, and compared against our post-processing method by our group (Sartorio et al., in prep.). Various on-the-fly models have been used by other groups before (e.g. Slavin2015; Martinez-Gonzalez2019; Vasiliev2024; Dedikov2025), however, their dust destruction models are not as complete as in our post-processing code, so that we chose to use Paperboats.
@@
  • We do not model the ejecta-dominated phase of the SNR evolution. Instead, we assume that the ejecta mass can always be neglected in comparison to the encountered ISM mass and insert only thermal energy and no ejecta mass. Thus, the SNR evolution starts immediately in the Sedov-Taylor phase instead of the ejecta-dominated phase. Furthermore, the pressure support of the ejecta is not modeled in this stage. The maximum amount of dust that would be encountered during the ejecta-dominated phase is 2–6% of the encountered mass after 10 kyr in our low ISM density simulations, depending on the considered ejecta mass (5–15 M). For our high and intermediate ISM density simulations, this fraction is $\sim$0.5–2.5% and thus not significant. At low densities, the ejecta mass would be significant only if the ejecta-dominated phase would be skipped entirely. Instead, we replace it by an immediate Sedov-Taylor phase, which has a similar evolution to the late ejecta-dominated phase. Thus, only a fraction of the 2–6% of the total dust mass is uncertain due to the missing ejecta mass and the dust destruction in our study can be expected to be only slightly underestimated in our low ISM density cases. However, since we start all of our simulations with an immediate Sedov-Taylor phase, concluded trends would not change qualitatively by considering ejecta mass.
@@
  • Because we do not include ejecta mass, we do not aim to resemble any specific SNR or SN type. Instead, we model SN type IIP and IIL in general by using a thermal energy of $10^{51}$ erg and having no pre-SN mass loss. Hence, our model only applies to this more general approach and no observed SN.
@@
  • Chemistry is not fully traced. With an assumed collisional ionization equilibrium and an electron-ion equilibration, cooling times, dust charges, and dust transport could be influenced. Non-equilibrium ionization could for example decrease the temperatures at the shock front (Slavin2015), resulting in less efficient thermal sputtering. However, tracing ionization states in detail, and ion and electron temperatures separately, is beyond the scope of this study.
@@
  • Although Arepo uses time step subcycling to handle unresolved cooling times $t_{\text{cool}}$ (Glover2007), the cooling length is not fully resolved during the whole simulation. The cooling length, $l_{\text{cool}}=c_{\text{s}}t_{\text{cool}}$ with the speed of sound $c_{\text{s}}$, can be as short as $3\times 10^{-7}$ pc in the shell of our high ISM density simulations because the ejecta cooling function peaks at $\sim$$2\times 10^{5}$ K. This results in undercooled gas at the shock front, overestimating gas temperatures and therefore thermal sputtering. This undercooling seems to be a general problem of SNR dust destruction studies due to the large difference of the cooling length and the used numerical resolution. Thus, the significance of the unresolved cooling length cannot be easily assessed.
@@
We emphasize that, while several assumptions may influence the absolute values of dust destruction in different ways, our model is applied consistently across all simulations. Therefore, the relative trends identified between simulations remain robust and meaningful.

### v2

@@ 3.4 Summary
To summarize, the SNR forward shock dust destruction depends on the properties of the surrounding medium (ISM density and turbulent Mach number), the dust grain species, and the considered destruction mechanisms. If only one of these parameters is varied, the dust destruction efficiency can change significantly. Generally, we find that the dust destruction efficiency decreases with lower ISM densities and higher turbulent Mach numbers. It is further reduced if carbonaceous instead of silicate dust is considered. Moreover, grain-grain collisions always generate many grains with sizes lower than 5 nm and can have a large impact on the destruction rate in intermediate- and high-density environments (up to 36%).
The lowest forward shock dust destruction in the first 10 kyr is found in the low-density ($n=1\,\mathrm{cm^{-3}}$) and strongly turbulent ($\mathcal{M}=3$) environment, where only 27% of carbonaceous dust is destroyed. With 0.85 M of carbonaceous dust, the least dust mass is destroyed in the $n=1\,\mathrm{cm^{-3}}$ and $\mathcal{M}=0.3$ case. However, the forward shock still has a high velocity of $v>250\,\mathrm{km\,s^{-1}}$ and is still hot at the end of these simulations, so that especially kinetic and thermal sputtering will process much more dust beyond the considered period. Compared to the average dust mass of 0.1–1 M that is observed to be produced by SNRs and that still needs to survive the high-velocity reverse shock, we conclude that SNRs are rather net dust destroyers, in particular if higher gas densities are considered.
@@ 4.1 Previous studies
Table 2: Comparison of this study to different forward shock dust destruction studies after 10 kyr.444[1] Dedikov and Vasiliev (2025), [2] Kirchschlager et al. (2024a), [3] Kirchschlager et al. (2022), [4] Martínez-González et al. (2019). Compared are the ISM inhomogeneity model, the average ISM density ($n$), the resolution, whether grain-grain collisions are considered or not ($g$), the initial grain-size distribution (GSD), and the destroyed dust mass (and rate) after the first 10 kyr of SNR evolution for $n\approx 1\,\mathrm{cm^{-3}}$ cases.
@@
Dedikov and Vasiliev (2025) conduct such a study, using a similar approach to ours by simulating the forward shock dust destruction over 100 kyr in a clumpy instead of turbulent medium, which is generated by fractal cubes. They use different inhomogeneity strengths and their densest clumps have roughly the same density contrasts between high- and low-density regions as our supersonic Mach number simulations. They do not consider magnetic fields nor grain-grain collisions, only kinetic and thermal sputtering. In their paper, they conclude that dense clumps can shield the dust efficiently, resulting in up to $\sim$20% less destruction over 100 kyr for an ISM with $n=1\,\mathrm{cm^{-3}}$ on average. After 10 kyr, they find 1.17 M ($\sim$39%) of destroyed dust when considering a homogeneous medium, decreasing with increasing inhomogeneity down to 0.67 M ($\sim$32%) of destroyed dust with the densest clumps so that these clumps already shield $\sim$7% of the dust at 10 kyr. Their considered grain species is not mentioned but the results match reasonably well to our carbonaceous (0.79–0.91 M or 25–33%) grain simulations with $n=1\,\mathrm{cm^{-3}}$ that include only sputtering, for which we also find significant filament shielding of 8% at 10 kyr. In an intermediate-density environment ($n=10\,\mathrm{cm^{-3}}$) with rather dense ISM clumps, $\sim$4.5 M ($\sim$50%) of dust is destroyed in the first 10 kyr, also fitting reasonably well to our results of the intermediate-density case with strong turbulence and carbonaceous grains (3.5–4.2 M or 46–63% of dust destroyed). This supports the conclusion that generally more dust is destroyed at higher ISM densities. At 100 kyr, however, this trend can reverse according to their study where 8.0–11.4 M dust is destroyed in the low-density case, but only 7.5 M in the inhomogeneous intermediate-density case.
@@
If our sputtering results for carbonaceous grains would also align quite well with the results of Dedikov and Vasiliev (2025) beyond 10 kyr, over 8 M of dust would be destroyed at 100 kyr, which is almost an order of magnitude more than what SNRs are observed to produce. Nevertheless, we note that the filaments in our study could behave differently than the clumpy medium of Dedikov and Vasiliev (2025) at lower shock velocities, and that we cannot estimate the dust evolution in our own simulations that include grain-grain collisions beyond 10 kyr based on this comparison, as this destruction mechanism could become more efficient at later SNR stages (see for example our $n=100\,\mathrm{cm^{-3}}$ case).
@@
Kirchschlager et al. (2024a) also investigate single SNRs in a turbulent ISM, similarly to us. However, they use SN-driven turbulence, which brings the caveats mentioned in Sect. 1 with it. After turbulence driving, they inject an SN in either a low- or intermediate-density area of $n=0.03\text{ or }0.7$ cm-3, respectively. Thus, although their average ISM density is always lower than our lowest density case ($n=1$ cm-3), the SN injection in their moderately dense region might be comparable to our simulations because their SNR does not encounter many voids, whereas we set off the SN in a void. Their dust properties and considered dust destruction mechanisms are similar to our study, but they only consider silicate grains.
@@
After 10 kyr, they find a destroyed dust mass of 2.8 M in the intermediate-density simulation, which is almost two times more than the destroyed silicate grain mass in our high Mach turbulence study (1.5 M). This disparity can come from multiple factors. Most importantly, the turbulence driving is different such that the filaments they produce appear much wider than ours. Furthermore, the average ISM density in the box is lower with the SN set off in a higher density environment. Thus, the dust destruction efficiency can change significantly (Kirchschlager et al. 2024a) and the encountered dust mass can be different. Moreover, their simulation is not fine-tuned to the first 10 kyr, meaning that the temporal and spatial resolution is rather coarse (250 yr and 0.5 pc) compared to our study (logarithmic time steps between 0.15–120 yr depending on the evolution time and 0.044 pc in the $n=1$ cm-3 cases), which can cause inaccuracies in the gas drag. With a lower spatial resolution and thus bigger grid cells, the dust velocity in a single cell is averaged over more grains, resulting in larger deviations compared to the real grain velocities. If the temporal resolution is too low, grains that are dragged with the forward shock may not be as efficiently decelerated by encountered ISM filaments as the gas so that the grains could be able to overtake the shock. This would leave the high-velocity grains exposed in a rather stationary ISM, leading to high relative gas-grain and grain-grain velocities.
@@
Thus, the higher destruction efficiency of Kirchschlager et al. (2024a) could be explained by an insufficient resolution. Other differences can add to these inaccuracies. First of all, the extrapolation from the 2D to the 3D dust mass is different and induces uncertainties. We multiply the averaged destruction efficiency of the three investigated 2D planes by the 3D SNR-encountered dust mass, whereas Kirchschlager et al. (2024a) fully extrapolate the dust mass from one 2D plane, using a correction factor they derive from comparing the SNRs in a homogeneous and an inhomogeneous medium. Thus, the real 3D destruction could be different. Secondly, they do not include partial vaporization, which was implemented in Paperboats by Kirchschlager et al. (2023). This slightly reduces the dust destruction for the largest grains in our simulations. Thirdly, the penetration depth of gas ions into dust grains was calculated considering oxygen ions in their study, but was updated in our study to consider hydrogen instead of oxygen ions, which have a different penetration depth. This value alters the size-dependent sputtering, so that significantly more small silicate grains are destroyed. Lastly, it should be mentioned that the dust destruction in a turbulent ISM also depends on the detailed ISM environment, as we show with the three different orthogonal planes in this paper. The differences in the destruction efficiencies could therefore also partly come from the different surrounding environment and positions of the filaments.
@@
Similarly to our study, however, the derived dust destruction efficiencies of Kirchschlager et al. (2024a) do not plateau after 10 kyr, but only start to plateau after 50 kyr in their study. Nevertheless, they find that some dust continues to be destroyed even when their simulations finish at 1 Myr, with a total of 37 M of destroyed dust, which is $\sim$13.5 times more than what they find at 10 kyr. If this trend is applied to our high Mach number and low-density ISM silicate simulation, it would mean that even in one of the lowest dust destruction cases, the SNR would destroy $1.5\,\mathrm{M_{\odot}}\times 13.5\approx 20\,\mathrm{M_{\odot}}$ of dust after 1 Myr, being 20 times higher than the observed dust production in SNRs. This implies, once again, that SNRs are rather net dust destroyers.
@@
Another useful comparison can be made to works that consider pre-SN feedback. Martínez-González et al. (2019) let a wind-driven bubble evolve in a homogeneous ISM with a density of 1 cm-3 before they inject the SN. The bubble’s shocked region is compressed and has a density that is roughly three to four orders of magnitude higher than the interior of the bubble in which the SN explodes, resulting in roughly the same difference between filaments and voids as in our low-density and supersonic turbulence simulations. It expands up to 23 pc in radius, meaning that it sweeps up $\sim$12.6 M of dust during the stellar wind evolution. They find that $\sim$0.36 M of dust (2.9% of the shell’s dust mass) is destroyed in the first 10 kyr after SN explosion if only thermal sputtering is considered and a mixture of silicate and carbonaceous grains is used. Although we also find that dust can be shielded by filaments, significantly more dust is shielded by their stellar wind shell than in our sputtering-only results with high Mach turbulence for either grain species (0.88–1.5 M and 25–45%). However, their bubble is larger and its shell is roughly 3–5 pc much thicker than our filaments, meaning that it can slow down the shock more efficiently. As a result, most of the dust in the shell has not yet been encountered after 10 kyr so that this dust cannot be destroyed. Furthermore, they do not include kinetic sputtering, which could be efficient and increase the destroyed dust fraction when the SNR forward shock starts to encounter the wind-driven bubble. At this point, grain-grain collisions could also become efficient due to the high dust densities and velocities, which would further increase the dust destruction.
Other studies that investigate the dust destruction of SNR forward shocks in a slightly inhomogeneous or turbulent ISM are Vasiliev and Shchekinov (2024) and Hu et al. (2019), respectively. The former uses the same code as Dedikov and Vasiliev (2025) but only introduces small density perturbations of 10% and does not compare their results to a homogeneous case so that the filament shielding efficiency cannot be assessed. Since they only study a monosized dust population, we cannot easily compare their study with ours because we use an MRN-distributed dust population, which leads to significantly different dust behavior. Hu et al. (2019) drive their turbulence with SNe, similar to Kirchschlager et al. (2024a). However, they study the lifetime of silicate and carbonaceous dust grains over a total of 400 Myr and several SNe. Thus, their destruction efficiencies are also not easily comparable to our study that only covers the first 10 kyr. However, similarly to us, they find that silicates are better destroyed than carbonaceous grains.
@@ 4.2 Caveats
  • The pre-SN effects of stellar winds and radiation were only mimicked simplistically in our study by setting off the SNe in voids because a full pre-SN feedback model is out of the scope of this paper. In particular the weakly turbulent and homogeneous runs did not incorporate realistic pre-SN structures. For a forthcoming model, we plan to improve this by allowing stellar winds and radiation to evolve naturally in separate simulations, and letting the SN explode inside of the wind-driven bubble.
@@
  • Although we consistently set off all SNe in voids, the detailed surrounding ISM structures are not the same, since the turbulence-driving results in different filamentary shapes for each simulation. This affects the 3D SNR evolution in the first 10 kyr and therefore also the dust destruction, resulting in an uncertainty, which could be quantified best by studying many different 3D ISM environments for every parameter set. Since the needed resources for this are too high, we assess this uncertainty with our investigation of the three orthogonal 2D planes, as discussed in Appendix D. We find a maximum uncertainty of 9.4% and include the standard deviation between those planes in our dust destruction results of Tables 1 and 4. This implies that the SNR dust destruction cannot be precisely predicted based on the initial parameters like ISM density and turbulent Mach number, but needs to take the exact SN ejection location (see also Appendix B) and the surrounding environment structures into account.
@@
  • Our post-processing approach does not allow the dust to feed heavy elements coming from destroyed dust grains back to the gas, which physically should happen and would induce for example a faster cooling of the gas. The exact impact of the dust feedback is so far unknown but might be significant. Since on-the-fly approaches are more expensive than post-processing simulations, we used the latter for our calculations to include a more extensive dust destruction model. However, this brings the drawbacks of no dust-to-gas feedback and coarser time steps, which gives inaccuracies in the destruction, transport, and charge processes that we tried to minimize by taking as many logarithmically spaced snapshots as computationally feasible. A comprehensive on-the-fly approach is currently being implemented, tested, and compared against our post-processing method by our group (Sartorio et al., in prep.). Various on-the-fly models were used by other groups before (e.g., Slavin et al. 2015; Martínez-González et al. 2019; Vasiliev and Shchekinov 2024; Dedikov and Vasiliev 2025); however, their dust destruction models were not as complete as in our post-processing code, so we chose Paperboats.
@@
  • We did not model the ejecta-dominated phase of the SNR evolution. Instead, we assumed that the ejecta mass can always be neglected in comparison to the encountered ISM mass and we inserted only thermal energy and no ejecta mass. Thus, the SNR evolution started immediately in the Sedov-Taylor phase instead of the ejecta-dominated phase. Furthermore, the pressure support of the ejecta was not modeled in this stage. The maximum amount of dust that would be encountered during the ejecta-dominated phase is 2–6% of the encountered mass after 10 kyr in our low ISM density simulations, depending on the considered ejecta mass (5–15 M). For our high- and intermediate-density ISM simulations, this fraction is $\sim$0.5–2.5% and thus not significant. At low densities, the ejecta mass would be significant only if the ejecta-dominated phase were skipped entirely. Instead, we replaced it by an immediate Sedov-Taylor phase, which has a similar evolution to the late ejecta-dominated phase. Thus, only a fraction of the 2–6% of the total dust mass is uncertain due to the missing ejecta mass and the dust destruction in our study can be expected to be only slightly underestimated in our low ISM density cases. However, since we started all of our simulations with an immediate Sedov-Taylor phase, concluded trends would not change qualitatively by considering ejecta mass.
@@
  • Because we did not include ejecta mass, we did not aim to resemble any specific SNR or SN type. Instead, we modeled SN type IIP and IIL in general by using a thermal energy of $10^{51}$ erg and having no pre-SN mass loss. Hence, our model only applies to this general approach and no observed SN.
@@
  • Chemistry was not fully traced. With an assumed collisional ionization equilibrium and an electron-ion equilibration, cooling times, dust charges, and dust transport could be influenced. Non-equilibrium ionization could for example decrease the temperatures at the shock front (Slavin et al. 2015), resulting in less efficient thermal sputtering. However, tracing ionization states in detail, and ion and electron temperatures separately, is beyond the scope of this study.
@@
  • Although Arepo uses time step subcycling to handle unresolved cooling times $t_{\text{cool}}$ (Glover and Mac Low 2007), the cooling length was not fully resolved during the whole simulation. The cooling length, $l_{\text{cool}}=c_{\text{s}}t_{\text{cool}}$ with the speed of sound $c_{\text{s}}$, can be as short as $3\times 10^{-7}$ pc in the shell of our high ISM density simulations because the ejecta cooling function peaks at $\sim$$2\times 10^{5}$ K. This results in undercooled gas at the shock front, overestimating gas temperatures and therefore thermal sputtering. This undercooling seems to be a general problem of SNR dust destruction studies due to the large difference of the cooling length and the used numerical resolution. Thus, the significance of the unresolved cooling length cannot be easily assessed.
@@
We emphasize that, while several assumptions may influence the absolute values of dust destruction in different ways, our model was applied consistently across all simulations. Therefore, the relative trends identified between simulations remain robust and meaningful.